Lynn wants to make 10kg of mixed nuts in which 1/4 of the nuts are almonds. If 1/3 of the Good Shell nuts and 1/5 of the Nutalicious nuts are almonds, how much of each type should be used in the mixture?

Answer by stanbon(75887)

Lynn wants to make 10kg of mixed nuts in which 1/4 of the nuts are almonds. If 1/3 of the Good Shell nuts and 1/5 of the Nutalicious nuts are almonds, how much of each type should be used in the mixture?
----
Almonds Equation: (1/3)g + (1/5)n = (1/4)10
Quantity Equation: g + n = 10 kg
-----
Modify::
20g + 12n = 15*10
20g + 20n = 20*10
------
Subtract and solve for "n"::
8n = 5*10
8n = 50
n = 6.25 kg (amt of Nutalicious nuts)
----
g = 10-6.25 = 3.75 kg (amt of Good Shell nuts)
===================
Cheers,
Stan H.